The world’s most powerful electromagnet is capable of producing a magnetic field with a strength of 45 T. If an electron enters this field at a right angle with a speed of 7.5 × 106 m/s, what will be the magnetic force acting on the electron?
1
Expert's answer
2019-05-06T10:45:02-0400

The magnetic force acting on the electron can be defined as follows


"\vec{F} = e (\vec{v} \times \vec{B})"

As long as the angle between the velocity and magnetic field vectors is 90 degree (and there is no other information on their directions), we can calculate the magnitude of the force (which is perpendicular to both aforementioned vectors):


"F = |e| v B = 1.6 \cdot 10^{-19} \cdot 7.5 \cdot 10^{6} \cdot 45 = 5.4 \cdot 10^{-11} \, N"

Answer: 5.4 x 10-11 N